OpenStudy (anonymous):

f(x) = x + 8 and g(x) = x2 – 6x – 7. Find f(g(2))

OpenStudy (hoblos):

7

OpenStudy (anonymous):

would it be -7?

hero (hero):

f(g(x)) = (x^2-6x-7) + 8 f(g(2)) = (2^2-6(2)-7) + 8 f(g(2)) = (4-12-7) + 8 f(g(2)) = 4-19 + 8 f(g(2))= 12-19 f(g(2)) = -7

OpenStudy (hoblos):

yeah yeah it is -7 sorryyyyyy